Applications of proportions.
Gus works twice as fast as sid. Together they can dig a garden in 4 hr. How long would it takes each person working alone.

Answer by stanbon(75887)

Gus works twice as fast as sid. Together they can dig a garden in 4 hr. How long would it takes each person working alone.
----
Sid time = 2x hr/job ; rate = 1/(2x) job/hr
-----
Gus time = x hr/job ; rate = 1/x job/hr
-----
Together time = 4 hr/job ; rate = 1/4 job/hr
-----
Equation:
rate + rate = together rate
1/(2x) + 1/x = 1/4
----------------
Multiply thru by 4x to get:
2 + 4 = x
x = 6 hrs (time for Gus to do the job)
2x = 12 hrs (time for Sid to do the job)
========================================
Cheers,
Stan H.
=================